# CodeIF-Bench: Evaluating Instruction-Following Capabilities of Large Language Models in Interactive Code Generation

### v3

@@ 5.1. RQ1: Performance in Static Conversation
Table 6. IA results in Dynamic Conversation. Values with darker colors indicate higher performance.
Level Model Turn 1-2 Turn 3-4 Turn 5-6 Turn 7-8 Turn 9-10 Turn 11-12 Turn 13-14 Turn 15-16 Turn 17-18 Avg.
L1 Claude-3.5-Sonnet 61.4 58.5 52.9 48.3 53.3 35.7 50.0 - - 51.5
GPT-4o 55.7 56.3 38.5 36.4 46.9 45.5 55.6 - - 47.8
DeepSeek-V3 51.4 54.5 50.9 36.2 35.3 50.0 70.0 - - 49.8
Qwen2.5-Coder-7B 35.0 34.8 16.1 17.0 5.7 3.7 0 - - 16.0
Qwen2.5-Coder-14B 40.0 36.2 19.5 13.9 20.8 21.1 18.2 - - 24.2
Qwen2.5-Coder-32B 46.4 39.9 25.0 12.1 11.5 9.5 8.8 - - 21.9
Qwen-Coder-Turbo 46.4 25.0 12.5 19.5 17.9 15.2 14.3 - - 21.5
L2 Claude-3.5-Sonnet 53.8 32.5 21.1 31.0 30.4 43.8 35.7 33.3 57.1 37.6
GPT-4o 16.2 8.8 10.3 2.8 0.0 0.0 0.0 0.0 0.0 4.2
DeepSeek-V3 50.0 31.2 19.7 23.3 22.0 22.2 12.5 18.2 25.0 24.9
Qwen2.5-Coder-7B 2.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.3
Qwen2.5-Coder-14B 12.5 6.2 6.2 2.6 0.0 5.3 0.0 2.9 0.0 4.0
Qwen2.5-Coder-32B 13.8 3.8 3.8 0.0 1.3 0.0 1.4 0.0 2.8 3.0
Qwen-Coder-Turbo 20.0 12.5 11.2 5.7 12.9 4.5 7.8 8.6 7.4 10.1
@@
Table 7. CA and IFE results in Dynamic Conversation. Values with darker colors indicate higher performance.
Level Model Turn 1-2 Turn 3-4 Turn 5-6 Turn 7-8 Turn 9-10 Turn 11-12 Turn 13-14 Turn 15-16 Turn 17-18 Avg.
CA IFE CA IFE CA IFE CA IFE CA IFE CA IFE CA IFE CA IFE CA IFE CA IFE
L1 Claude-3.5-Sonnet 63.6 56.1 73.5 57.9 70.0 50.9 68.5 46.0 61.7 40.9 58.1 37.8 51.0 32.9 - - - - 63.8 46.1
GPT-4o 60.7 58.2 69.9 55.4 62.0 45.2 59.7 41.4 45.4 31.2 32.7 24.0 36.6 27.9 - - - - 52.4 40.5
DeepSeek-V3 56.4 51.8 65.0 49.4 67.4 47.9 50.7 32.6 45.7 30.1 48.0 29.3 59.6 35.1 - - - - 56.1 39.5
Qwen2.5-Coder-7B 39.6 39.3 46.5 35.3 31.6 22.8 23.8 16.1 18.1 12.0 11.7 7.8 7.0 4.0 - - - - 25.5 19.6
Qwen2.5-Coder-14B 43.6 39.6 53.9 41.1 45.8 31.9 33.4 21.0 26.4 16.4 23.4 14.7 10.4 8.4 - - - - 33.8 24.7
Qwen2.5-Coder-32B 51.8 48.6 52.8 42.1 43.1 30.8 28.0 17.9 19.7 12.5 12.0 7.5 10.4 6.3 - - - - 31.1 23.7
Qwen-Coder-Turbo 51.4 47.5 48.2 36.6 28.2 18.9 26.2 16.4 24.2 15.4 21.6 13.4 8.9 5.9 - - - - 29.8 22.0
L2 Claude-3.5-Sonnet 59.4 56.2 52.5 41.4 50.7 36.8 44.7 30.7 39.4 26.7 38.8 26.1 37.8 25.9 33.9 23.0 49.6 35.3 45.2 33.6
GPT-4o 16.9 15.6 14.3 11.9 13.8 10.0 8.9 5.9 5.8 3.8 2.1 1.2 1.4 0.8 1.2 0.7 0.4 0.2 7.2 5.6
DeepSeek-V3 54.4 51.2 49.5 38.7 44.7 32.2 40.3 26.4 38.9 24.6 34.0 20.7 30.6 18.1 18.3 10.9 19.5 11.8 36.7 26.1
Qwen2.5-Coder-7B 3.8 3.8 1.0 0.7 0.7 0.5 0.6 0.3 0.5 0.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.7 0.6
Qwen2.5-Coder-14B 13.8 13.8 6.8 5.2 7.0 5.0 2.4 1.5 0.9 0.5 7.8 4.4 5.6 3.0 2.8 1.5 0.0 0.0 5.2 3.9
Qwen2.5-Coder-32B 15.0 13.8 7.0 5.6 5.5 4.0 2.4 1.6 0.7 0.4 0.8 0.5 1.0 0.6 0.9 0.5 1.4 0.8 3.9 3.1
Qwen-Coder-Turbo 22.5 22.5 15.9 12.6 16.8 11.4 8.0 5.0 12.2 7.7 9.4 5.8 8.2 5.0 8.0 4.7 5.6 3.3 11.8 8.7
5.2. RQ2: Performance in Dynamic Conversation
@@
Table 6 presents the Instruction Accuracy (IA) results on CodeIF-Bench under the Dynamic Conversation. The introduction of feedback (compiler feedback and unfollowed instructions) significantly increases dialogue rounds. In the L-1 task, LLMs show different adaptability. Stronger models, such as Claude, demonstrate better robustness, and small-parameter open-source models gradually lose their ability to follow current-round instructions as the number of rounds increases. This indicates that stronger LLMs can effectively utilize external feedback to correct code and reduce the negative impact of the number of rounds for following current instructions. However, in the L-2 dataset, the IA performance of most LLMs shows varying degrees of decline. Notably, GPT-4o exhibits significant performance degradation compared to its performance in L-1. This indicates that the long context may prevent the LLM from effectively utilizing feedback and benefiting from it.
Table 7 presents the Conversation Accuracy (CA) results. Same as Static Conversation, we exclude “Functionality Extension” instructions in this experiment. In the L-1 task, compared to Static Conversation without feedback, LLMs like Claude and GPT4o showed substantial performance improvements (Comparison of Avg.: 44.5 vs 63.8 and 42.5 vs 52.4, respectively). However, the Qwen2.5-Coder series displayed only marginal improvements or even declines, revealing that the conversation performance gap between LLMs in dialogue tasks also stems from their differential ability to utilize feedback for code improvement. Interestingly, while DeepSeek-V3’s average CA score surpassed GPT4o’s, its IFE was lower, indicating that DeepSeek-V3 can effectively use feedback to enhance instruction compliance but requires more interaction rounds, thereby reducing interaction efficiency. This observation validates that our IFE and CA metrics provide a comprehensive evaluation of LLM performance. In L-2 data, Claude and DeepSeek-V3 outperformed other models, with CA metrics showing significant improvement over static dialogue settings, further demonstrating that improving models’ feedback utilization represents a promising direction for enhancing multi-round dialogue performance, especially when longer task context is required.
@@
Figure 4. IFR results in Dynamic Conversation.
Additionally, we investigated the Instruction Forgetting Rate (IFR) of the LLMs. As shown in Figure 4, in the L1 task, the IFR of most LLMs gradually increases as the number of dialogue rounds grows, with Claude’s IFR surging to over 40% in rounds 13–14. In the L2 dataset, as dialogue rounds increase, Claude and DeepSeek-V3, which perform well in CA—show a gradually rising IFR, further suggesting that longer context length is a key factor in LLM forgetting. Notably, Qwen2.5-Coder-7B and 32B models exhibit a distinct pattern where IFR first rises and then declines. This behavior is because weaker LLMs are less capable of following new instructions, and in subsequent rounds, most of the instructions received are repeated prompts for previously correct but forgotten responses. This phenomenon mirrors real-world scenarios where users prompt LLMs to ”recall” previous states, highlighting the critical role of user feedback and guidance in conversational performance.
Based on the above analysis, we identify that LLMs are more inclined to complete the current instruction and struggle to determine the scope of code edits. When faced with multiple instructions in extended conversations, LLMs fail to precisely delineate the editing boundaries for each instruction. This leads to this failure mode: when attempting to complete the current instruction, the LLM accidentally edits the previously correct part of the code, the model appears to “forget”. Consequently, Claude and DeepSeek-V3, which have stronger IA capabilities, exhibit higher IFR, thereby affecting CA, while other LLMs with weaker IA capabilities received more reminders from users about forgetting them because they removed more instructions that they could not perform correctly, resulting in an initial increase followed by a decrease in IFR.
RQ2 Summary: In Dynamic Conversation with explicit external feedback, the LLM’s ability to effectively utilize feedback can substantially mitigate the negative effects of extended contexts on instruction following, particularly in long-context scenarios like repository-level code generation. Moreover, LLM struggles to determine the appropriate edit boundaries to accidentally edit the previously correct part of the code, which may result in instruction forgetting, ultimately degrading its instruction-following performance during interactions.
Table 8. IA and CA results across 8 interaction turns. Darker colors indicate higher performance.
Model Turn-1 Turn-2 Turn-3 Turn-4 Turn-5 Turn-6 Turn-7 Turn-8 Avg.
IA CA IA CA IA CA IA CA IA CA IA CA IA CA IA CA IA CA
Deepseek-Coder-6.7B 37.1 37.1 64.3 42.9 70.0 30.0 51.4 14.3 47.1 10.0 70.6 4.3 59.2 11.9 46.2 55.7 21.6
Deepseek-Coder-6.7B-CoT 34.3 34.3 61.4 41.4 71.4 24.3 55.7 20.0 47.1 7.1 63.6 1.4 54.7 13.4 44.7 54.1 20.4
OpenCodeInterpreter-DS-6.7B 32.9 32.9 58.6 40.0 64.3 32.9 44.3 20.0 40.0 11.4 61.7 7.1 47.3 14.9 39.3 48.6 22.8
CodeLlama-7B 35.7 27.1 47.1 30.0 64.3 15.7 52.9 8.6 44.3 4.3 83.8 0.0 64.8 0.0 20.1 50.2 12.4
CodeLlama-7B-CoT 34.3 22.9 12.9 7.1 20.0 0.0 15.7 2.9 25.7 0.0 21.7 0.0 10.2 0.0 8.2 17.3 4.8
OpenCodeInterpreter-CL-7B 30.0 30.1 22.9 10.0 67.1 20.0 27.1 5.7 27.1 1.4 17.4 0.0 25.9 1.4 15.3 28.1 9.9
CodeLlama-13B 30.0 30.0 38.6 25.7 44.3 12.9 45.7 5.7 50.0 7.1 52.2 0.0 40.5 0.0 23.3 40.7 11.8
CodeLlama-13B-CoT 21.4 21.4 17.1 7.1 55.7 7.1 44.3 0.0 35.7 0.0 44.5 0.0 56.0 0.0 12.7 36.1 5.2
OpenCodeInterpreter-CL-13B 22.9 22.9 37.1 22.9 62.9 18.6 30.0 8.6 41.4 7.1 20.2 5.7 47.4 6.0 27.6 36.3 13.2
5.3. RQ3: Performance on Various VI Strategies
@@
Figure 5. The IA results on various VI strategies.
To evaluate LLMs’ instruction following performance across different instruction types, we implemented a 2-round experimental design: the first round comprised basic programming tasks, while the second round focused on verifiable instructions. This design effectively controlled for potential confounding effects from varying dialogue round counts. As illustrated in Figure 5, the L-1 dataset revealed significant performance variations in IA scores across instruction types, particularly in “Annotation Coverage”, “Code Standard”, and “Functionality Extension”. We also find that the underperforming Turns 5, 7, and 8 in Table 4 mostly correspond to these three instruction strategies. This suggests that LLMs may have insufficient training data for non-functional requirements (e.g., annotations and code style) and functional enhancement scenarios, resulting in comparatively weaker performance.
@@
5.4. RQ4: Exploration of IF Enhancement
To investigate this question, we selected two representative methods: Chain-of-Thought (Kojima et al., 2023) (CoT), a prompt engineering approach, and OpenCodeInterpreter (Zheng et al., 2024), a fine-tuning-based method. For CoT, we adopted the original prompt template from its paper and applied it consistently across each dialogue round. OpenCodeInterpreter employs LLMs as simulators to interactively generate dialogue training data for code generation tasks, thereby enhancing the model’s performance in interactive coding scenarios. Given its substantial training dataset of 68K samples and generated by GPT-4, we chose it as the representative fine-tuning-based approach. For comparison, we evaluated the baseline model of OpenCodeInterpreter alongside DeepSeek-Coder-6.7B-Instruct, CodeLlama-7B-Instruct, and CodeLlama-13B-Instruct.
As shown in Table 8, CoT demonstrates a negative impact on improving LLMs’ instruction-following performance, with all metrics (including IA and CA) underperforming relative to the baseline. Our findings indicate that even when LLMs engage in structured, step-by-step reasoning, they frequently fail to produce correct outputs, and even the thought process is affected in its original ability to follow instructions, suggesting that easy prompt engineering alone is insufficient to enhance instruction following ability and may need carefully designed templates. Furthermore, while fine-tuning methods yield modest improvements in CA for DeepSeek-Coder-6.7B-Instruct and CodeLlama-13B-Instruct, the gains are marginal, and IA exhibits varying degrees of degradation. This may be because OpenCodeInterpreter only uses LLM to generate data and does not further optimise data quality, thereby introducing noisy data caused by such as LLM hallucinations. This implies that although fine-tuning shows promise for improving LLM instruction-following in dialogue settings, further research is necessary to optimize its efficacy, particularly in areas such as data quality.
RQ4 Summary: Simply relying on thinking prompt templates (e.g., Chain-of-Thought) is insufficient to enhance an LLM’s instruction-following ability in conversational settings. And such reasoning processes may even degrade its baseline performance. While fine-tuning-based methods present a promising direction for improvement, further research is required to optimize their effectiveness.
6. Threats to Validity

### v4

@@ 5.2 RQ2: Performance in Dynamic Conversation
Table 6: IA results in Dynamic Conversation. Values with darker colors indicate higher performance
Level Model Turn 1 Turn 2 Turn 3 Turn 4 Turn 5 Turn 6 Turn 7 Turn 8 Turn 9 Turn 10 Turn 11 Avg.
L1 Claude-3.5-Sonnet 55.7 71.4 78.8 66.1 60.0 57.1 56.5 50.0 66.7 - - 68.2
GPT-4o 57.1 54.3 64.2 55.7 58.8 58.1 42.9 60.0 50.0 - - 59.0
DeepSeek-V3 55.7 60.0 56.5 64.4 59.2 53.3 65.2 53.8 63.6 - 60.2
Qwen2.5-Coder-7B 48.6 41.4 44.1 39.3 35.2 36.4 27.6 26.9 25.0 - - 43.9
Qwen2.5-Coder-14B 45.7 51.4 52.9 45.9 45.8 50.0 37.5 44.4 50.0 - - 50.0
Qwen2.5-Coder-32B 55.7 52.9 51.5 54.1 47.1 39.4 35.0 62.5 35.7 - - 52.3
L2 Claude-3.5-Sonnet 52.5 45.0 42.5 40.0 28.9 14.7 29.0 25.0 26.3 33.3 27.3 37.2
GPT-4o 47.5 32.5 40.0 25.0 28.9 21.6 42.9 35.7 27.3 23.8 22.2 33.9
DeepSeek-V3 50.0 47.5 35.0 35.0 28.9 43.8 33.3 54.5 42.1 58.8 38.5 41.8
Qwen2.5-Coder-7B 35.0 10.0 20.0 20.0 10.0 17.9 11.1 14.3 20.0 18.8 19.4 18.8
Qwen2.5-Coder-14B 47.5 32.5 25.0 25.0 17.9 14.3 23.5 14.8 25.9 25.9 20.8 27.3
Qwen2.5-Coder-32B 40.0 27.5 17.5 12.5 12.8 0.0 2.7 5.7 2.9 0.0 0.0 14.3
@@
Table 7: CA and CIF results in Dynamic Conversation. Values with darker colors indicate higher performance
Level Model Turn 1 Turn 2 Turn 3 Turn 4 Turn 5 Turn 6 Turn 7 Turn 8 Turn 9 Turn 10 CIF
L1 Claude-3.5-Sonnet 48.6 58.7 66.5 65.8 64.6 58.8 50.4 35.7 - - 74.5
GPT-4o 49.0 55.3 60.7 59.8 61.3 57.1 35.5 44.8 - - 71.9
DeepSeek-V3 48.6 58.0 63.8 62.3 65.1 55.2 52.0 30.8 - - 72.3
Qwen2.5-Coder-7B 44.8 48.0 50.3 50.0 41.7 35.4 19.3 22.8 - - 52.3
Qwen2.5-Coder-14B 46.1 54.2 58.3 57.2 54.2 52.1 38.8 32.3 - - 65.0
Qwen2.5-Coder-32B 47.2 56.9 60.1 59.6 60.4 51.4 36.1 36.4 - - 65.9
L2 Claude-3.5-Sonnet 40.0 46.5 44.3 46.8 43.7 42.1 38.1 32.1 24.7 22.0 47.8
GPT-4o 42.3 40.7 40.7 41.5 36.6 37.0 34.6 30.4 19.5 22.9 40.0
DeepSeek-V3 41.2 47.5 48.0 48.7 44.7 45.3 40.4 35.5 33.2 35.3 54.0
Qwen2.5-Coder-7B 31.2 23.0 22.2 19.7 15.8 13.3 12.5 7.4 11.7 7.2 18.0
Qwen2.5-Coder-14B 35.5 32.7 32.0 27.7 29.2 22.9 23.5 12.2 14.4 14.1 31.3
Qwen2.5-Coder-32B 34.7 30.5 23.5 17.8 15.1 9.2 7.0 5.7 4.1 2.4 13.8
@@
Table 6 reports the Instruction Accuracy (IA) results under the Dynamic Conversation setting on the CodeIF-Bench dataset. Due to the requirement to incorporate feedback from the base task $I_{1}$, an additional round of interaction is required compared to Static Conversation. In L-1 tasks, most stronger LLMs exhibit strong adaptability, maintaining stable performance across interaction rounds. However, weaker LLM models such as the Qwen2.5-Coder series continue to exhibit unstable performance. This indicates that stronger LLMs can effectively utilize external feedback to correct code and reduce the negative impact of the number of rounds for following current instructions. However, in the L-2 dataset, the IA performance of most LLMs shows varying degrees of decline. For instance, Claude-3.5-Sonnet—one of the strongest closed-source models—achieved a score of merely 14.7 at Turn 6. This indicates that the long context may prevent the LLM from effectively utilizing feedback and benefiting from it.
Table 7 presents the Conversation Accuracy (CA) results. Same as Static Conversation, we exclude “Functionality Extension” instructions in this experiment. In the L-1 task, similar to the Static Conversation setting, all LLMs exhibit a trend of initially rising and then declining CA, indicating that even with feedback, their instruction-following ability at the session level gradually weakens as conversations progress. Compared with the feedback-free Static Conversation results (final-round CA), models such as Claude and DeepSeek-V3 show notable improvements (68.0 → 74.5 and 57.4 → 72.3, respectively). This demonstrates that incorporating feedback effectively enhances LLMs’ instruction-following capability during interactive dialogues. Furthermore, the comparison between DeepSeek-V3 and Claude (74.5 vs. 72.3) suggests that feedback allows relatively weaker models to achieve performance comparable to stronger ones, thereby reducing potential deployment costs. In the L-2 task, CIF also achieves substantial gains over the Static Conversation setting without feedback (e.g., Claude: 25.8 → 47.8; DeepSeek-V3: 25.6 → 40.0), further highlighting feedback’s role in strengthening instruction following. However, CA still declines notably as conversations lengthen, indicating that LLMs continue to face challenges in maintaining consistent instruction following across extended multi-turn interactions.
We further analysed the Instruction Forgetting Rate (IFR), as illustrated in Figure 4. Across both L-1 and L-2 tasks, IFR progressively increases with dialogue depth, suggesting that LLMs struggle to maintain conversational state when handling multiple instructions simultaneously. During such exchanges, models often lose control over editing operations, inadvertently modifying previously correct content and exhibiting forgetting behaviour. Interestingly, in both L-1 and L-2 settings, a subset of IFR trajectories shows an initial rise followed by a decline, corresponding to dialogue rounds where users restate previously forgotten instructions. This finding indicates that enabling LLMs to recognise and recover forgotten or previously executed instructions can potentially improve their instruction-following performance in dynamic interaction scenarios.
@@
Figure 4: IFR results in Dynamic Conversation
RQ2 Summary: In the Dynamic Conversation setting, stronger LLMs effectively leverage feedback to maintain stable instruction-following performance in L-1 tasks. However, in L-2 tasks, extended context severely hinders feedback utilization, leading to declines in both IA and CA. Although feedback improves instruction following numbers during conversations, performance still degrades over longer interactions as Instruction Forgetting Rate (IFR) increases. These results highlight that enabling LLMs to recognise and recover forgotten or previously executed instructions is a potential factor for enhancing robustness in long-context, multi-turn scenarios.
5.3 RQ3: Performance on Various VI Strategies
@@
Figure 5: The IA results on various VI strategies
To evaluate LLMs’ instruction following performance across different instruction types, we implemented a 2-round experimental design: the first round comprised basic programming tasks, while the second round focused on verifiable instructions. This design aims to controlle for potential confounding effects from varying dialogue round counts. As illustrated in Figure 5, the L-1 dataset revealed significant performance variations in IA scores across instruction types, particularly in “Annotation Coverage”, “Code Standard”, and “Functionality Extension”. We also find that the underperforming Turns 5, 7, and 8 in Table 4 mostly correspond to these three instruction strategies. This suggests that LLMs may have insufficient training data for non-functional requirements (e.g., annotations and code style) and functional enhancement scenarios, resulting in comparatively weaker performance.
@@
5.4 RQ4: Exploration of IF Enhancement
To further enhance LLMs’ instruction following performance in interactive code generation, one intuitive and lightweight approach is to improve the prompting strategy. Unlike model fine-tuning or reinforcement optimization, which require extensive data and computational resources, prompt enhancement focuses on optimizing the interaction layer between the developer and the model, making it more practical. Therefore, we explore whether carefully designed prompting strategies can improve LLMs’ ability to interpret and execute evolving developer instructions across multiple turns. To investigate this question, we examine three representative prompting strategies that capture different aspects of context management and reasoning:
  • Full History (FH): preserves all prior user–LLM interactions as input for the next round (consistent with Static Conversation). This retains all prior user–LLM interactions as input for the next round, enabling the model to leverage complete conversational context.
  • Chain-of-Thought (CoT)[cot]: instructs the LLM to explicitly generate a reasoning process before producing the final answer in each round. The objective is to investigate whether explicit reasoning enhances the model’s instruction following performance in multi-turn scenarios.
  • Cumulative Instruction (CI): retains only the user instructions from all previous rounds, excluding model responses, as context for subsequent turns. For example, in $i$-th round, compared to FH’s prompt $(I_{1},A_{1},I_{2},A_{2},...,I_{i})$, the prompt is $(I_{1},I_{2},...,I_{i})$. This approach highlights the trajectory of developer intent while mitigating the accumulation of errors introduced by earlier model outputs, allowing assessing whether selective context preservation improves overall performance.
The Instruction Accuracy (IA) results of the three prompt-enhancing methods are presented in Table 8. As shown in the results, in the L-1 task, consistent with single-turn performance, the CoT method improves adherence to current instructions through explicit reasoning, whereas CI does not produce significant improvement. In contrast, in the L-2 task, CI achieves the best performance, while CoT remains effective but suboptimal. These findings indicate that accumulated conversational context can degrade LLM performance, while CoT mitigates this degradation through structured reasoning. However, retaining only essential user instructions (as in CI) further enhances multi-turn instruction following.
Table 8: The IA results for different prompting methods. The backbone model is DeepSeek-V3
L1 FH 60.0 80.0 84.3 77.4 55.7 71.8 29.1 54.9 - - 64.2
CoT 57.1 78.6 84.3 74.3 60.0 70.4 43.5 53.6 - - 65.2
CI 58.6 81.4 85.7 72.9 55.7 77.7 24.8 58.3 - - 64.4
L2 FH 50.0 55.0 62.5 52.5 27.5 30.0 2.5 45.5 45.5 27.5 39.8
CoT 55.0 62.5 60.0 55.0 37.5 37.5 10.0 55.0 47.5 35.0 45.5
CI 52.5 62.5 57.5 70.0 47.5 27.5 5.0 70.0 55.0 40.0 48.8
Table 9: The CA results for different prompting methods. The backbone model is DeepSeek-V3. Tokens represents the average total number of tokens required per conversation
Level Model Turn-1 Turn-2 Turn-3 Turn-4 Turn-5 Turn-6 Turn-7 Turn-8 Turn-9 Avg. Tokens.
L1 FH 60.0 77.6 78.0 77.4 72.8 62.4 57.4 - - 69.4 8.2K
CoT 57.1 75.4 76.4 75.4 72.0 65.7 63.1 - - 69.3 21.7K
CI 58.6 78.5 78.6 78.3 73.3 71.9 67.7 - - 72.4 4.2K
L2 FH 50.0 51.2 54.2 51.2 35.0 32.1 29.3 31.6 33.6 40.9 71.3K
CoT 55.0 57.5 55.8 55.7 46.0 42.5 37.1 41.3 41.1 48.0 79.1K
CI 52.5 57.5 59.2 60.6 54.0 50.4 45.4 46.9 47.8 52.7 54.3K
Table 9 presents the Conversational Accuracy (CA) results as well as the computational costs (Tokens). These results further demonstrate that CI not only achieves the highest CA in both L-1 and L-2 tasks but also incurs the lowest computational overhead. This underscores the promise of context compression—extracting key elements from dialogue history—to maintain both efficiency and instruction following. Future work could explore adaptive context selection mechanisms that dynamically identify and preserve critical historical information, thereby enabling LLMs to enhance conversational performance.
RQ4 Summary: Excessive conversational history can hinder instruction following, but reasoning-based prompting (CoT) and cumulative instruction (CI) alleviate this issue from different perspectives. CI, in particular, offers a more efficient and scalable approach for sustaining LLM performance across multi-turn interactions. Exploring to preserve critical historical information is a potential method for efficiently enhancing LLMs’ conversational performance.
5.5 Case Study
Figure 6: Case analysis of DeepSeek-V3 in L-2 Tasks
To provide a more intuitive illustration of LLM performance on CodeIF-Bench, Figure 6 presents representative examples of DeepSeek-V3’s responses in L-2 tasks. In the left-hand example, during the 7-th interaction round, the model was instructed to use a class member variable within the given context. While it successfully identified and referenced the variable, it mistakenly invoked a non-existent method. This suggests that although LLMs can recognize relevant contextual information, their ability to apply it accurately remains limited. In real-world development scenarios, LLMs must not only retrieve relevant repository context but also comprehend and apply it correctly to follow user instructions—highlighting a key challenge in repository-level code generation. The right-hand example shows that, as conversational context accumulates over multiple rounds, the model erroneously modifies a previously correct line of code in the sixth round. This demonstrates that cumulative dialogue context can degrade an LLM’s instruction-following ability. Hence, enabling LLMs to extract effective information from accumulated dialogue and recognize forgotten or previously executed instructions, represents a possible direction for enhancing their robustness in multi-turn interactions.
In summary, within real-world repository development, LLMs must effectively manage two intertwined forms of context—repository context and conversational context. The ability to integrate and reason over both is crucial for improving instruction-following performance.
6 Threats to Validity